<Problem 2>

Answer the following questions about the square matrix A of order 3:

A= 3 01 1 2 -1 -2-2 1

III. The square matrix B of order 3 is diagonalizable and meets AB=BA.

prove that any eigenvector p of A is also an eigenvector of B.

IV. Find the square matrix B of order 3 that meets B2 = A, where B is

diagonalizable and all eigenvalues of B are positive.

V. The square matrix X of order 3 is diagonalizable and meets AX = XA.

When tr(AX) = d, find the maximum of det(AX) as a function of d.

Here, d is positive real and all eigenvalues of X are positive. In addition,

tr(M) is the trace (the sum of the main diagonal elements) of the square matrix M,

and det(M) is the determinant of the matrix M.
Answer #1

Given, 1 2 0 1 A Here de tA) =6 and eigen values of A are: 1,2,3. So A is diagonalizable and each eigen space corresponding t(1) To find B sit ² A First we have to find eigen vectors of A. For Ar! A-1769) - 7 23)-(0) /201 -1.1- 1-2- 2012 2017030 ni-let Then A= PD pt where p= diag (1,2,3) het B = PVD p . where ro= on il O OV3 Then p² = (PVD pt 72 - pppy A so required B= PlA = 120 Given -21 . .) - P diag (1,2,3) p! Ax=XA , To Qx)=0 = x must be of the form P Do PT for some diagonal matrin Di= diag> det(x) = 2 So manimum value of de+(Axis d3 when na 24=32 = 0/3 ie when (ny, z) = (as de ida).